OpenStudy (anonymous):

Nancy has a sewing project which calls for panels which use 1 1/4 yd of fabric each. She has 4 1/2 yd of fabric. How many panels can she cut from the fabric? A. 2 B. 3 C. 4 D. 5

sam (.sam.):

= (4 1/2)/(1 1/4) = (9/2)/(5/4) = 9/2(4/5) or 36/10 = 18/5 or 3 3/5 Answer: 3 panels